OpenStudy (anonymous):

the displacement from equilibrium of an object of harmonic motion on the end of a spring is y=1/3cos12t-1/4sin12t where y is measure in feet and t is the time in seconds. Determine the position and velocity of the object when t=pi/8.

OpenStudy (anonymous):

Determine the position and velocity at t=pi/8 Displacement and Position mean the same thing. so plug in t-pi/8 into the equation y=1/3cos12t-1/4sin12t the derivative of y=1/3cos12t-1/4sin12t with respect to t is velocity. so take the derivative, plug in t=pi/8, and you get velocity